The electric field in a microwave traveling through air has amplitude 0.50 mV/m and frequency 30 GHz. Find the amplitude and frequency of the magnetic field.
Answer #1

As

For electromagnetic Wave

Relation between peak value of electric and peak value of magnetic field is given by

E0/B0=c

So

Amplitude of magnetic field=A=E0/c=(0.5010-3)/(3108)=1.67 10-12 T

Frequency of magnetic field will remain same and=30 GHz